# Gradient Existence and Energy Finiteness of Local Minimizers in the Wasserstein $L^{\infty}$ Topology for Binary-Star Systems

### v1

@@ 2.1 Notations and problem setting
In general, it is easier to consider the case that the density is centered at 0, which is in ${R}_{0}\left(\mathbb{R}^{3}\right)$.
In subsection 2.2 one can see when the topology satisfies certain conditions, if the $z$-component $J=J_{z}$ of the angular momentum $\boldsymbol{J}$ is fixed, the problem of minimizing $E(\rho,v)$ locally in $R_{0}(\mathbb{R}^{3})\times V(\mathbb{R}^{3})$ is equivalent to a minimization problem of the energy $E_{J}(\rho)$ locally in $R_{0}(\mathbb{R}^{3})$. Here $E_{J}(\rho)$ corresponds to uniform rotation with the angular momentum $\boldsymbol{J}=J\hat{e}_{z}=(0,0,J)^{T}$ specified a priori, and $E_{J}(\rho)$ is defined as the following:
@@
Due to this uniform rotation observation, in the presence of 2-body systems, although the systems rotate in universe, after fixing a time, we can still assume that they fall within two disjoint regions $\Omega_{m}$ and $\Omega_{1-m}$, widely separated relative to $\frac{J^{2}}{\mu^{2}}$, where $\mu=m(1-m)$ is their reduced mass. Note that the same symbol $\mu$ is also used as for Lebesgue measure, provided no ambiguity arises. For the planet’s mass $m\in(0,1)$ and the star’s mass (1m)(1-m), we consider $E_{J}(\rho)$ is minimized subject to the constraint
$W_{m,J}\coloneq\left\{\rho(m)=\rho_{m}+\rho_{1-m}\in R\left(\mathbb{R}^{3}\right)\mid\int\rho_{m}=m,\text{spt }\rho_{m}\subset\Omega_{m},\text{spt }\rho_{1-m}\subset\Omega_{1-m}\right\}$ (17)
@@
Fix two points $y_{m}$ and $y_{1-m}$ in $\mathbb{R}^{3}$ from the plane $z=0$, which are separated by $\eta=\frac{J^{2}}{\mu^{2}}$, i.e. $\eta=\left|y_{m}-y_{1-m}\right|$. The $\Omega_{m}$ and $\Omega_{1-m}$ are defined as closed balls in $\mathbb{R}^{3}$ centered at $y_{m}$ and $y_{1-m}$, whose size and separation scale with $\eta$ as the following:
@@
The distance separating $\Omega_{m}$ and $\Omega_{1-m}$, and the diameter of their union is given by:
@@
The reason we set the separation $\eta=\frac{J^{2}}{\mu^{2}}$ in the definitions above is inspired by the Kepler problem. Given two point masses $m$ and 1m1-m, rotating with angular momentum $J>0$ about their fixed center of mass. Assume their separation is $d$, the gravitational energy plus kinetic energy is $\frac{\mu}{d}+\frac{J^{2}}{2\mu d^{2}}$, which reaches its minimum at separation $d=\eta$.
@@
We hope that the radii of $\Omega_{m}$ and $\Omega_{1-m}$ increase as $\eta$ increases is to guarantee the supports of constrained minimizers (stars or planets) will fall in the interior of $\Omega_{m}\cup\Omega_{1-m}$. This is because McCann shows the size of stars or planets will not expand too much as $\eta$ increases. As discussed in [31, Section 6], it lays the groundwork for discussion on the constrained minimizers being local minimizers (here and in what follows, unless otherwise specified, “local” will always refer to “local under the topology induced by the Wasserstein $L^{\infty}$ distance”). Therefore, the minimizers are solutions to (EP’) due to Theorem 2.25.
When the distance between two objects is large, the gravitational interaction between them becomes very weak. As a result, the stars are not significantly deflected, and their supports remain limited. Therefore, they can be approximated as point masses, which simplifies the analysis.
When $\eta$ is large (which occurs when $\mu$ is small or $J$ is large), since the gravitational interaction between two bodies seems small, the stars are not significantly deflected, and their sizes are negligible compared to the distance between them. Hence, one can expect a stable, slowly rotating equilibrium to exist in which fluid components with masses $m$ and 1m1-m lie near $y_{m}$ and $y_{1-m}$. It is indeed true as McCann discussed in [31, Section 6]. For prescribed angular momentum, we also know the energy $E(\rho,v)$ is bounded from below on ${R}_{0}\left(\mathbb{R}^{3}\right)$ by the non-rotating energy $e_{0}$. However, as in Morgan [32], McCann [31, Example 3.6] demonstrates that this bound — although approached — will not be attained, see also [30]. Therefore, instead of searching for a global energy minimizer in binary-star system, one is forced to settle for local minimizers of $E(\rho,v)$ in an appropriate topology.
@@
In fact, the velocity fields $v$ may be topologized in any way which makes ${V}\left(\mathbb{R}^{3}\right)$ a topological vector space, since the velocity distribution minimizing the kinetic energy is always given by uniform rotation [31, Section 3] [23, Lemma 2.2.2] as we will describe in subsection 2.2. However, the choice of topology for ${R}_{0}\left(\mathbb{R}^{3}\right)$ is quite delicate: McCann pointed out local energy minimizer will not exist if the topology of $R_{0}(\mathbb{R}^{3})$ is inherited from a topological vector space ([31, Remark 3.7]), i.e. ${R}_{0}\left(\mathbb{R}^{3}\right)$ is a subset of a topological vector space (see [33, Chapter 1, Section 16]). While we partially acknowledge this perspective, we argue that the focus should be on $R(\mathbb{R}^{3})$ rather than $R_{0}(\mathbb{R}^{3})$. For further details, one can refer to Proposition 5.15 and Remark 5.19 in subsection 5.2. We will further compare this case with the case that the topology induced by the Wasserstein $L^{\infty}$ distance, to highlight the advantages of considering topology induced by the Wasserstein $L^{\infty}$ distance in such problems.
@@
Let $\rho\geq 0$, $\sigma\geq 0$ be the density functions in $\mathbb{R}^{3}$ with mass $\int\rho=m_{1}<\infty,\int\sigma=m_{2}<\infty$. $\bar{x}(\rho)$ and $\bar{x}(\sigma)$ denote the centers of mass, $I(\rho)$ and $I(\sigma)$ denote the moments of inertia, function $r$ is given in (2.1).
@@
The (1) case is obvious, since $m_{1}+m_{2}=0$ implies $\rho+\sigma$ is 0 almost everywhere. For (2) case, by definition $\left(m_{1}+m_{2}\right)\bar{x}(\rho+\sigma)=\int(\rho+\sigma)x=m_{1}\bar{x}(\rho)+m_{2}\bar{x}(\sigma)$, then
$I(\rho+\sigma)$ $=\int(\rho+\sigma)r^{2}(x-\bar{x}(\rho+\sigma))$
$=\int\rho r^{2}(x-\bar{x}(\rho)+\bar{x}(\rho)-\bar{x}(\rho+\sigma))+\int\sigma r^{2}(x-\bar{x}(\sigma)+\bar{x}(\sigma)-\bar{x}(\rho+\sigma))$
@@
$\int\rho r^{2}(x-\bar{x}(\rho)+\bar{x}(\rho)-\bar{x}(\rho+\sigma))=$ $\int\rho r^{2}(x-\bar{x}(\rho))+\int\rho r^{2}(\bar{x}(\rho)-\bar{x}(\rho+\sigma))$
$+2\int\rho\langle x-\bar{x}(\rho),\bar{x}(\rho)-\bar{x}(\rho+\sigma)\rangle_{2}$
Notice by definition of center of mass, we have $\int\rho\langle x,\bar{x}(\rho)-\bar{x}(\rho+\sigma)\rangle_{2}=m_{1}\langle\bar{x}(\rho),\bar{x}(\rho)-\bar{x}(\rho+\sigma)\rangle_{2}=\int\rho\langle\bar{x}(\rho),\bar{x}(\rho)-\bar{x}(\rho+\sigma)\rangle_{2}$, thus $2\int\rho\langle x-\bar{x}(\rho),\bar{x}(\rho)-\bar{x}(\rho+\sigma)\rangle_{2}=0$, therefore,
$\int\rho r^{2}(x-\bar{x}(\rho)+\bar{x}(\rho)-\bar{x}(\rho+\sigma))$ $=I(\rho)+m_{1}r^{2}\left(\frac{m_{2}}{m_{1}+m_{2}}(\bar{x}(\rho)-\bar{x}(\sigma))\right)$
@@
$\int\sigma r^{2}(x-\bar{x}(\sigma)+\bar{x}(\sigma)-\bar{x}(\rho+\sigma))=I(\sigma)+\frac{m_{2}\cdot m_{1}^{2}}{\left(m_{1}+m_{2}\right)^{2}}r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))$
@@
Physically it’s natural to assume those mass, center of mass and moment of inertia finite. However, one may notice as long as the mass $m_{1}$ and $m_{2}$ are finite and one of the centers of mass is finite, the result can still hold even though the values may be infinite.
@@ 2.2 Uniform rotation minimizes kinetic energy
This subsection reviews several results of McCann [31], Jang and Seok [23]. First, we introduce that the problem of minimizing $E(\rho,v)$ is equivalent to a minimization problem of the energy $E_{J}(\rho)$ with respect to uniform rotation, that is, the fluid rotates uniformly about its center of mass. Then we show that with uniform rotation setting, the conservation of mass $\partial_{t}\rho+\nabla(\rho v)=0$ automatically holds true and then the Euler-Poisson equations (EP) can be simplified to (EP’) as Jang and Seok describes [23].
When we define the total energy $E(\rho,v)$, we can notice that the velocity fields $v$ only appears in the kinetic energy $T(\rho,v)$. Therefore, we can first can consider the velocity minimizer of $T(\rho,v)$ with a given density $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$. McCann points out that $T(\rho,v)$ is uniquely minimized by a uniform rotation:
@@
Given $\rho$ in ${R}_{0}\left(\mathbb{R}^{3}\right)$ fixed, we define the Hilbert space $H\coloneq L^{2}\left({\mathbb{R}^{3},d\rho(x)}\right)\subset V\left(\mathbb{R}^{3}\right)$, where $d\rho(x)=\rho(x)d^{3}x$, with inner product $v,w>_{H}\coloneq\int(v\cdot w)\rho$, in particular $v,v>_{H}=2T(\rho,v)$. One can check: (1) $\hat{e_{z}}\times x\in H$; (2) if $v\in{V}\left(\mathbb{R}^{3}\right)$ satisfies $T(\rho,v)<\infty$ and the constraint ${J}_{z}(\rho,v)=J$, then it is in the space $G=\left\{v\in H\middle|<v,~\hat{e_{z}}\times x>_{H}=J\right\}$; (3) if $v$ is in $G$, then $v$ can be expressed as $v=v_{g}+v_{o}$, where $v_{g}=\frac{J}{I(\rho)}\left(\hat{e_{z}}\times x\right)=\omega\left(\hat{e_{z}}\times x\right)$ and $v_{o}$ is orthogonal to $\hat{e_{z}}\times x$ in $H$, that is $v_{0},\hat{e_{z}}\times x>_{H}=0$. Therefore,
@@
We have the following corollary:
@@
Let $J\geq 0$ and $\rho\in{R}\left(\mathbb{R}^{3}\right)$, and define $\omega:=\frac{J}{I(\rho)}$. Let $V(\mathbb{R}^{3})$ be a topological vector space, and the topology on ${R}_{0}\left(\mathbb{R}^{3}\right)$ is such that the map taking $\rho$ to $I(\rho)$ is continuous. Then $(\rho,v)$ minimizes $E(\rho,v)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ subject to the constraint $J_{z}(\rho,v)=J$ if and only if $\rho$ minimizes $E_{J}(\rho)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)$ and $v(x)=\omega\hat{e}_{z}\times x$. Here $E_{J}(\rho)=U(\rho)-\frac{G(\rho,\rho)}{2}+T_{J}(\rho)$.
@@
By Lemma 4.7 (v), we know that ${R}_{0}\left(\mathbb{R}^{3}\right)$ is endowed with the topology induced by the Wasserstein $L^{\infty}$ distance, then it satisfies the conditions required in the above corollary, namely, that the mapping $\rho\mapsto I(\rho)$ is continuous.
@@
Corollary 2.18 implies even if $(\rho,v)$ minimizes $E(\rho,v)$ only locally on ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ subject to the constraint $J_{z}(\rho,v)=J$, we still have $v(x)=\omega\hat{e}_{z}\times x$, which, to some extent, strengthens the result of Proposition 2.17. Actually it comes from McCann’s result [31, Corollary 3.2], where the minimum of $T(\rho,\widehat{v}(t))$ or $E(\rho,\widehat{v}(t))$ along the curve $\widehat{v}(t)=(1-t)\tilde{v}(x)+t\omega\hat{e}_{z}\times x$ is considered.
Remark 2.21.
In the arguments above, we only consider minimization on ${R}_{0}\left(\mathbb{R}^{3}\right)$ or ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$. By translation invariant of energy we can see the results also holds over ${R}_{a}\left(\mathbb{R}^{3}\right)$ or ${R}_{a}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ for any $a\in\mathbb{R}^{3}$, and furthermore, it can actually be extended to minimization over ${R}\left(\mathbb{R}^{3}\right)$ or ${R}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$. However, since the solutions to (EP’) considered later are in $W_{m,J}$, in particular, they have compact support, hence the above result suffices for our purposes.
@@
where $\omega=\frac{J}{I\left(\rho\right)}$, then $({\rho}(t,x),v(t,x))$ gives a uniform rotating binary star system (or star-planet system) such that $E(\rho,v)$ is local minimum.
Now we have understood one motivations of defining $E_{J}$. Actually, $E_{J}$ can be derived through another approach, inspired by Jang and Seok [23].
@@
Where $\nabla_{y}$ means the gradient operator w.r.t. $y$, and $\nabla_{x}$ means the gradient operator with w.r.t. $\widetilde{x}$, where $\widetilde{x}=R_{-\omega t}y$ above.
@@
Where $\nabla$ means the gradient operator w.r.t. $x$ by default. We call (EP’) reduced Euler-Poisson equation. One can also reverse this process to demonstrate that (EP’) can imply (EP) as well.
If we further assume $P^{\prime}$ and $\nabla\widetilde{\rho}$ exists, then (EP’) becomes
@@
such that $\text{spt}\left(\widetilde{\rho}_{i}\right)$ is connected for every $i\in 1,\ldots,N$ and mutually disjoint. Notice when $s>0$, $A^{\prime\prime}(s)$ exists and $A^{\prime\prime}(s)s=P^{\prime}(s)$ where $A$ is given in (1). Then in the region $\{\widetilde{\rho}>0\}$, dividing (24) by $\widetilde{\rho}$, one has
@@
Someone might be a bit confused when they first see the integration definition of $A(\rho)$ in (1). But it turns out with this definition we can have $A^{\prime}(\rho)\rho-A(\rho)=P(\rho)$. If $P(\rho)$ is continuous differentiable, we can have $A^{\prime\prime}(\rho)\rho=P^{\prime}(\rho)$ where $\rho>0$, which has been used above to obtain (EL0’).
Remark 2.23.
One can notice in (23) we assume the density rotates uniformly around 0, which implicitly assumes that the center of mass is located at the origin. However, from a physical standpoint, it is more appropriate to consider the object rotating about its center of mass. Also, from a mathematical standpoint, thanks to the similar arguments in Proposition 2.17 and Corollary 2.18 above, and Theorem 2.25, we know that if $(\widetilde{\rho},v)$ is the minimizer, then the fluid should indeed rotate around its center of mass. Luckily, when we consider Euler-Lagrange equation (EL), although the corresponding minimizer does not necessarily have $(0,0,0)^{T}$ as its center of mass, we can apply similar arguments as above. Therefore, instead of (EL0’) sometimes it is more convenient for us to first consider the following kind of Euler-Lagrange equation:

### v2

@@ 2.1 Notations and problem setting
Here, the support of $\rho$, denoted by spt $\rho$, is the smallest closed set carrying the full mass of $\rho$. In general, it is easier to consider the case that the density is centered at 0, which is in ${R}_{0}\left(\mathbb{R}^{3}\right)$.
We also consider uniform rotation with the angular momentum $\boldsymbol{J}=J\hat{e}_{z}=(0,0,J)^{T}$ specified a priori. In this case, the energy is given by $E_{J}(\rho)$ as follows:
@@
Due to this uniform rotation observation, after fixing a time, we can assume that the components of the 2-body system fall within two disjoint regions $\Omega_{m}$ and $\Omega_{1-m}$, widely separated relative to $\frac{J^{2}}{\mu_{r}^{2}}$, where $\mu_{r}=m(1-m)$ is their reduced mass. For the planet’s mass $m\in(0,1)$ and the star’s mass (1m)(1-m), we consider $E_{J}(\rho)$ to be minimized subject to the constraint
$W_{m,J}\coloneq\left\{\rho(m)=\rho_{m}+\rho_{1-m}\in R\left(\mathbb{R}^{3}\right)\mid\int_{\mathbb{R}^{3}}\rho_{m}\,dx=m,\text{spt }\rho_{m}\subset\Omega_{m},\text{spt }\rho_{1-m}\subset\Omega_{1-m}\right\}$ (18)
@@
Fix two points $y_{m}$ and $y_{1-m}$ in $\mathbb{R}^{3}$ from the plane $z=0$, which are separated by $\eta=\frac{J^{2}}{\mu_{r}^{2}}$, i.e. $\eta=\left|y_{m}-y_{1-m}\right|$. The $\Omega_{m}$ and $\Omega_{1-m}$ are defined as closed balls in $\mathbb{R}^{3}$ centered at $y_{m}$ and $y_{1-m}$, whose size and separation scale with $\eta$ as the following:
@@
The distance separating $\Omega_{m}$ and $\Omega_{1-m}$, and the diameter of their union are given by:
@@
The reason we set the separation $\eta=\frac{J^{2}}{\mu_{r}^{2}}$ in the definitions above is inspired by the Kepler problem. Given two point masses $m$ and 1m1-m, rotating with angular momentum $J>0$ about their fixed center of mass, if we assume their separation is $d$, then the gravitational energy plus the kinetic energy is $\frac{\mu_{r}}{d}+\frac{J^{2}}{2\mu_{r}d^{2}}$, which reaches its minimum at separation $d=\eta$. When $\eta$ is large (which occurs when $\mu_{r}$ is small or $J$ is large), since the gravitational interaction between two bodies seems small, one can expect a stable, slowly rotating equilibrium to exist in which fluid components with masses $m$ and 1m1-m lie near $y_{m}$ and $y_{1-m}$. It is indeed true as McCann discussed in [31, Section 6].
@@
We choose that the radii of $\Omega_{m}$ and $\Omega_{1-m}$ increase as $\eta$ increases in order to guarantee the supports of constrained minimizers (stars or planets) will fall in the interior of $\Omega_{m}\cup\Omega_{1-m}$. This is because McCann shows the size of stars or planets will not expand too much as $\eta$ increases. As discussed in [31, Section 6], it lays the groundwork for discussion on the constrained minimizers being local minimizers (here and in what follows, unless otherwise specified, “local” will always refer to “local under the topology induced by the Wasserstein $L^{\infty}$ distance”). Therefore, the minimizers are solutions to (EP’) due to Theorem 2.22.
For prescribed angular momentum, we also know the energy $E(\rho,v)$ is bounded from below on ${R}_{0}\left(\mathbb{R}^{3}\right)$ by the non-rotating energy $e_{0}$. However, as in Morgan [32], McCann [31, Example 3.6] demonstrates that this bound — although approached — will not be attained (see also [30]). We clarify this phenomenon in the following remark:
Remark 2.14.
(i) Non-attainability: By the definition of $E_{J}(\rho)$, we know $\forall\rho\in{R}\left(\mathbb{R}^{3}\right),E_{J}(\rho)\geq E_{0}(\rho)\geq e_{0}(1)=e_{0}$. But $e_{0}$ cannot be attained by $E_{J}(\rho)$ if $J>0$. This is because $E_{J}(\rho)=E_{0}(\rho)+T_{J}(\rho)$. If $E_{0}(\rho)=e_{0}$, then $\rho$ is a minimizer of $E_{0}$ and hence has compact support [31, Theorem 3.5 (v)]. It makes $I(\rho)>0$ thus $T_{J}(\rho)>0$, $E_{J}(\rho)>e_{0}$; if $E_{0}(\rho)>e_{0}$, then, of course, $E_{J}(\rho)>e_{0}$.
(ii) Approachability: It turns out $e_{0}$ is a strictly concave function of the mass $m$ [31, Theorem 3.5 (ii)]. Hence, one can construct $\rho\in R_{0}\left({R}^{3}\right)$ such that $E_{J}(\rho)-e_{0}<\epsilon$ where $\epsilon$ is arbitrarily small, and the test function can be chosen as $\rho(x)=\sigma_{m}(x)+\sigma_{1-m}(x-y)$ with $|y|$ sufficiently large, where $\sigma_{m}$ and $\sigma_{1-m}$ denote the non-rotating minimizers with mass $m$ and 1m1-m respectively. The idea comes from Lieb, Morgan, and McCann [28, 30, 32, 31].
Therefore, instead of searching for a global energy minimizer in binary-star system, one is forced to settle for local minimizers of $E(\rho,v)$ in an appropriate topology.
@@
In fact, the velocity field $v$ may be topologized in any way which makes ${V}\left(\mathbb{R}^{3}\right)$ a topological vector space, since the velocity distribution minimizing the kinetic energy is always given by uniform rotation [31, Section 3] [23, Lemma 2.2.2] as we will describe in subsection 2.2. However, the choice of topology for ${R}_{0}\left(\mathbb{R}^{3}\right)$ is quite delicate: McCann pointed out local energy minimizer will not exist if the topology of $R_{0}(\mathbb{R}^{3})$ is inherited from a topological vector space ([31, Remark 3.7]), i.e. ${R}_{0}\left(\mathbb{R}^{3}\right)$ is a subset of a topological vector space (see [33, Chapter 1, Section 16]). While we partially acknowledge this perspective, we argue that the focus should be on $R(\mathbb{R}^{3})$ rather than on $R_{0}(\mathbb{R}^{3})$. For further details, one can refer to Proposition 5.14 and Remark 5.17 in subsection 5.2. We will further compare this case with the case that the topology induced by the Wasserstein $L^{\infty}$ distance, to highlight the advantages of considering the topology induced by the Wasserstein $L^{\infty}$ distance in such problems.
@@
Let $\rho\geq 0$, $\sigma\geq 0$ be the density functions in $\mathbb{R}^{3}$ with mass $\int_{\mathbb{R}^{3}}\rho\,dx=m_{1}<\infty,\int_{\mathbb{R}^{3}}\sigma\,dx=m_{2}<\infty$. $\bar{x}(\rho)$ and $\bar{x}(\sigma)$ denote the centers of mass, $I(\rho)$ and $I(\sigma)$ denote the moments of inertia, function $r$ is given in (2.1).
@@
The case (1) is obvious, since $m_{1}+m_{2}=0$ implies $\rho+\sigma$ is 0 almost everywhere. For the case (2), by definition $\left(m_{1}+m_{2}\right)\bar{x}(\rho+\sigma)=\int_{\mathbb{R}^{3}}(\rho+\sigma)x\,dx=m_{1}\bar{x}(\rho)+m_{2}\bar{x}(\sigma)$, then
$I(\rho+\sigma)$ $=\int_{\mathbb{R}^{3}}(\rho+\sigma)r^{2}(x-\bar{x}(\rho+\sigma))\,dx$
$=\int_{\mathbb{R}^{3}}\rho r^{2}(x-\bar{x}(\rho)+\bar{x}(\rho)-\bar{x}(\rho+\sigma))\,dx+\int_{\mathbb{R}^{3}}\sigma r^{2}(x-\bar{x}(\sigma)+\bar{x}(\sigma)-\bar{x}(\rho+\sigma))\,dx$
@@
$\int_{\mathbb{R}^{3}}\rho r^{2}(x-\bar{x}(\rho)+\bar{x}(\rho)-\bar{x}(\rho+\sigma))\,dx=$ $\int_{\mathbb{R}^{3}}\rho r^{2}(x-\bar{x}(\rho))\,dx+\int_{\mathbb{R}^{3}}\rho r^{2}(\bar{x}(\rho)-\bar{x}(\rho+\sigma))\,dx$
$+2\int_{\mathbb{R}^{3}}\rho\langle x-\bar{x}(\rho),\bar{x}(\rho)-\bar{x}(\rho+\sigma)\rangle_{2}\,dx$
Notice by the definition of the center of mass, we have
$\int_{\mathbb{R}^{3}}\rho\langle x,\bar{x}(\rho)-\bar{x}(\rho+\sigma)\rangle_{2}\,dx=m_{1}\langle\bar{x}(\rho),\bar{x}(\rho)-\bar{x}(\rho+\sigma)\rangle_{2}=\int_{\mathbb{R}^{3}}\rho\langle\bar{x}(\rho),\bar{x}(\rho)-\bar{x}(\rho+\sigma)\rangle_{2}\,dx$
Thus $2\int_{\mathbb{R}^{3}}\rho\langle x-\bar{x}(\rho),\bar{x}(\rho)-\bar{x}(\rho+\sigma)\rangle_{2}\,dx=0$, therefore,
$\int_{\mathbb{R}^{3}}\rho r^{2}(x-\bar{x}(\rho)+\bar{x}(\rho)-\bar{x}(\rho+\sigma))\,dx$ $=I(\rho)+m_{1}r^{2}\left(\frac{m_{2}}{m_{1}+m_{2}}(\bar{x}(\rho)-\bar{x}(\sigma))\right)$
@@
$\int_{\mathbb{R}^{3}}\sigma r^{2}(x-\bar{x}(\sigma)+\bar{x}(\sigma)-\bar{x}(\rho+\sigma))\,dx=I(\sigma)+\frac{m_{2}\cdot m_{1}^{2}}{\left(m_{1}+m_{2}\right)^{2}}r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))$
@@
Physically it’s natural to assume those mass, center of mass, and moment of inertia are finite. However, one may notice as long as the mass $m_{1}$ and $m_{2}$ are finite and one of the centers of mass is finite, the result can still hold even though the values may be infinite.
@@ 2.2 Uniform rotation minimizes kinetic energy
This subsection reviews several results of McCann [31], Jang and Seok [23]. First, we mention that the problem of minimizing $E(\rho,v)$ is equivalent to a minimization problem of the energy $E_{J}(\rho)$ with respect to uniform rotation, that is, the fluid rotates uniformly about its center of mass. Then we show that with uniform rotation setting, the conservation of mass $\partial_{t}\rho+\nabla(\rho v)=0$ automatically holds true, and then the Euler-Poisson equations (EP) can be simplified to (EP’) as Jang and Seok describe [23].
When we define the total energy $E(\rho,v)$, we can notice that the velocity field $v$ only appears in the kinetic energy $T(\rho,v)$. Therefore, we can first consider the velocity minimizer of $T(\rho,v)$ with a given density $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$. McCann points out that $T(\rho,v)$ is uniquely minimized by a uniform rotation:
@@
Given $\rho$ in ${R}_{0}\left(\mathbb{R}^{3}\right)$ fixed, we define the Hilbert space $H\coloneq L^{2}\left({\mathbb{R}^{3},d\rho(x)}\right)\subset V\left(\mathbb{R}^{3}\right)$, where $d\rho(x)=\rho(x)dx$, with inner product $v,w>_{H}\coloneq\int_{\mathbb{R}^{3}}(v\cdot w)\rho\,dx$, in particular $v,v>_{H}=2T(\rho,v)$. One can check: (1) $\hat{e_{z}}\times x\in H$; (2) if $v\in{V}\left(\mathbb{R}^{3}\right)$ satisfies $T(\rho,v)<\infty$ and the constraint ${J}_{z}(\rho,v)=J$, then it is in the space $G=\left\{v\in H\middle|<v,~\hat{e_{z}}\times x>_{H}=J\right\}$; (3) if $v$ is in $G$, then $v$ can be expressed as $v=v_{g}+v_{o}$, where $v_{g}=\frac{J}{I(\rho)}\left(\hat{e_{z}}\times x\right)=\omega\left(\hat{e_{z}}\times x\right)$ and $v_{o}$ is orthogonal to $\hat{e_{z}}\times x$ in $H$, that is $v_{0},\hat{e_{z}}\times x>_{H}=0$. Therefore,
@@
In fact, McCann obtained the following corollary, which, to some extent, strengthens Proposition 2.17.
@@
Let $J\geq 0$ and $\rho\in{R}\left(\mathbb{R}^{3}\right)$, and define $\omega:=\frac{J}{I(\rho)}$. Let $V(\mathbb{R}^{3})$ be a topological vector space, and let the topology on ${R}_{0}\left(\mathbb{R}^{3}\right)$ be such that the map taking $\rho$ to $I(\rho)$ is continuous. Then $(\rho,v)$ minimizes $E(\rho,v)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ subject to the constraint $J_{z}(\rho,v)=J$ if and only if $\rho$ minimizes $E_{J}(\rho)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)$ and $v(x)=\omega\hat{e}_{z}\times x$. Here $E_{J}(\rho)=U(\rho)-\frac{G(\rho,\rho)}{2}+T_{J}(\rho)$.
@@
By Lemma 4.5 (v), we know if ${R}_{0}\left(\mathbb{R}^{3}\right)$ is endowed with the topology induced by the Wasserstein $L^{\infty}$ distance, then the topology satisfies the conditions required in the above corollary, namely, that the mapping $\rho\mapsto I(\rho)$ is continuous.
@@
In the arguments above, we only consider minimization on ${R}_{0}\left(\mathbb{R}^{3}\right)$ or ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$. By translation invariance of energy we can see the results also hold over ${R}_{a}\left(\mathbb{R}^{3}\right)$ or ${R}_{a}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ for any $a\in\mathbb{R}^{3}$, and furthermore, it can actually be extended to minimization over ${R}\left(\mathbb{R}^{3}\right)$ or ${R}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$.
@@
where $\omega=\frac{J}{I\left(\rho\right)}$, then $({\rho}(t,x),v(t,x))$ gives a uniform rotating binary star system (or star-planet system) such that $E(\rho,v)$ is the local minimum.
Now we have understood one motivation of defining $E_{J}$. Actually, $E_{J}$ can be derived through another approach, inspired by Jang and Seok [23].
@@
where $\nabla_{y}$ means the gradient operator w.r.t. $y$, and $\nabla_{\widetilde{x}}$ means the gradient operator w.r.t. $\widetilde{x}$, where $\widetilde{x}=R_{-\omega t}y$ above.
@@
where $\nabla$ means the gradient operator w.r.t. $x$ by default. We call (EP’) the reduced Euler-Poisson equation. One can also reverse this process to demonstrate that (EP’) can imply (EP) as well.
If we further assume $P^{\prime}$ and $\nabla\widetilde{\rho}$ exist, then (EP’) becomes
@@
such that $\text{spt}\left(\widetilde{\rho}_{i}\right)$ is connected for every $i\in 1,\ldots,N$ and the supports are mutually disjoint. Notice when $s>0$, $A^{\prime\prime}(s)$ exists and $A^{\prime\prime}(s)s=P^{\prime}(s)$ where $A$ is given in (1). Then in the region $\{\widetilde{\rho}>0\}$, dividing (25) by $\widetilde{\rho}$, one has
@@
Sometimes, instead of (EL0’), it is more convenient for us to first consider the following kind of Euler-Lagrange equation: